Lactic acid: 0.300 L
0.500 mol/L
Ka: 1.4×10^-4
Sodium lactate: 0.300 L
0.300 L

a. Calculate pH of the lactic acid before it is mixed with NaLac to form buffer
b. Calculate the pH of NaLac before it is mixed with HLac to form buffer
c. Calculate pH of the buffer
d. What will the pH of the above solution be after 5 cm3 of 0.200 mol/dm3 NaOH has been added to only 100cm3 of this buffer
1
Expert's answer
2020-09-29T06:49:08-0400

a) H+ = under root Concentration * dissociation constant .


So , H+ = Under root 0.5 * 1.4 * 10-4 .


H+ = Under root 0.7 * 10-4 . = 0.83 * 10-2 .


So , PH = 1.91 answer .


( b )

concentration of NaC = 0.5 mol / liter . = C .


PH = 7 + PK /2 + log C / 2 .


PH = 7 + 3.85 / 2 + log 0.5 / 2 .


PH = 7 + 1.925 - 0.3 / 2


PH = 7 + 1.925 - 0.15 = 8.775 answer .

( c )

PH = Pk + log concentration of salt / concentration of acid .


PH = 3.85 + log 0.5 / 0.5 = 3.85 . answer .


d)


After adding NaOH PH will increases .


After adding NaOH this react with acid and form salt in which amount of acid decreases and amount of salt increases .